# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 7.6 Advocating for 3-way Exchanges and NDD-Initiated Chains
To operationalize NEPKE, it was necessary to establish a web-based data entry system. This system would enable each transplant center to enter their patient and donor information directly. The process of setting up this system and addressing bureaucratic requirements for NEPKE spanned about a year. During this period, Ünver developed the software for NEPKE using a priority matching algorithm formulated in Roth et al. (2005b). [^77] When only 2-way exchanges are feasible, a priority matching also maximizes the number of patients who receive a transplant (Roth et al., 2005b). Even though Delmonico initially indicated that only 2-way kidney exchanges were feasible in New England, for our research purposes, Ünver also coded priority matching algorithms that allowed for larger size exchanges.
After NEPKE became operational in 2006, kidney exchanges in New England were determined by Ünver’s software for a few years. Every few months, our team received anonymized patient-donor data from NEPKE, and we ran the software to derive any new kidney exchanges. This was an exciting period for us, as we were directly impacting and saving lives with our ideas. Hence, upon receiving anonymized data from NEPKE, Ünver and I would immediately inspect it to estimate how many transplants the software would determine.
In one of these instances, the software produced one more transplant than we thought would be feasible when 3-way exchanges were allowed in the system. This observation came as a big surprise. At that time, our intuition for the maximal-size kidney exchange was derived from Roth et al. (2005b), admittedly in a model that only permitted 2-way exchanges. To address this limitation, we would then utilize some of the technical details in our proofs to estimate the maximum number of transplants if larger-size exchanges were also allowed. This unexpected discovery revealed that our intuition was evidently incomplete.
Subsequently, we conducted a detailed analysis of the impact of allowing larger-size kidney exchanges on the number of transplants. Based on this analysis, we discovered that, due to the partial order structure induced by the blood type donation system:
@@
    In a sufficiently large patient-donor pool, the maximum number of transplants can be achieved if 3-way and 4-way kidney exchanges are allowed, and
@@
    Much of these gains over 2-way exchanges are from 3-way exchanges. [^78] For small pools with 25 patient-donor pairs, the gain was approximately a 30% increase in transplants from kidney exchanges. In large pools with 100 patient-donor pairs, this increase reduced to around a 20% additional transplants from kidney exchanges.
@@
Given these discoveries, and despite Delmonico’s initial reservations, we managed to convince our New England partners to include 3-way kidney exchanges in the system, publishing our findings in Saidman et al. (2006) and Roth et al. (2007). [^79] With 3-way exchanges allowed in our system, a priority matching no longer guaranteed the maximization of the number of transplants in the system. Consequently, we adopted an alternative algorithm that maximized the number of transplants à la Roth et al. (2007).
@@
Our collaboration with New England partners also led to another important invention. Before this collaboration, altruistic donors were not a major source of living donor kidneys in the U.S. According to the UNOS database, only 8 such transplants were conducted in the U.S. before 2000 (OPTN, 2024a). However, this number was increasing in the early 2000s, offering the possibility to initiate donor chains with an altruistic non-directed donor (NDD), rather than a patient-donor pair volunteering for list exchange. Moreover, since these NDD-chains do not interfere with the DD lists, they were not subject to the same ethical concerns that hindered the utilization of list exchanges in NEPKE.
@@
NDD chains also presented another possibility. As mentioned earlier, in their Consensus Statement, Abecassis et al. (2000) emphasized that all procedures in kidney exchange should be performed simultaneously, primarily to prevent a situation where the donor of a pair donates a kidney as part of a donor exchange, but their intended patient cannot receive a transplant later on due to the other donor backing out or becoming unavailable. However, this concern was significantly reduced with an NDD chain that starts with an altruistic donation. Since all patients receive a transplant before their paired donor donates to the next patient in the chain, the worst that could happen is the chain remaining shorter than it could be. [^80] Over time, even this worst-case scenario proved to be rare and of negligible consequence. Using data from 344 non-simultaneous NDD chains implemented by the National Kidney Registry from 2008 through May 2016, which generated 1,568 transplants, only 20 of these chains were broken due to reneging (Cowan et al., 2017). Interestingly, the average size of the broken chains was 4.8, compared to 4.5 for unbroken chains. Based on these observations, we formulated non-simultaneous NDD chains in Roth et al. (2006). While this invention was not pursued by NEPKE, it was later adopted by two other kidney exchange programs: first by the Alliance for Paired Donation (APD), [^81] Starting in 2006, our team also established a partnership with APD. Based on Roth et al. (2006, 2007), APD included 3-way and 4-way exchanges in their allocation mechanism from its inception and incorporated non-simultaneous NDD chains shortly after. Not only did Ünver introduce the concept of non-simultaneous NDD chains to Mike Rees, the head of APD, but the software used by APD from 2006 to 2008—just like the code of NEPKE—was also developed by him. See Anderson et al. (2015) for details of our subsequent partnership with APD, later renamed Alliance for Paired Kidney Donation (APKD). launched in 2006 in the Midwest, and shortly thereafter by the National Kidney Registry, launched in 2008.
@@
Due to its logistical advantages, the role of non-simultaneous NDD chains in kidney exchange programs in the U.S. increased considerably over time, particularly due to the National Kidney Registry’s success in enrolling altruistic non-directed donors into its system (Cowan et al., 2017; Agarwal et al., 2019).
@@ 7.7 Incentivized Kidney Exchange
While our pioneering role in the evolution and widespread adoption of kidney exchange programs worldwide stands as one of our most consequential policy achievements to date, a key element is still lacking in a vast majority of these programs—one that has the potential to significantly enhance their efficiency.
@@
As we briefly discussed in Section 7.5, partly due to events that led to the foundation of NEPKE—a program that became a blueprint for the rest of the world—participation in kidney exchange programs remains largely restricted to patients who are biologically incompatible with their donors. The efficiency loss resulting from this restriction is substantial. For the case of the U.S., Sönmez et al. (2020) estimates that including blood-type compatible pairs in kidney exchange could increase the number of kidney exchange transplants by as much as 160%.
@@
A handful of policies have been proposed and promoted by various groups to encourage the participation of compatible pairs in kidney exchange programs. One such policy, incentivized kidney exchange (IKE), was developed by Ünver and me in the early 2010s, as detailed in Sönmez and Ünver (2015) and later partially published in Sönmez et al. (2020). However, this policy has yet to be adopted in practice.
IKE refers to a nuanced strategy to encourage blood-type compatible pairs to participate in kidney exchange. The primary idea involves incentivizing such pairs by offering the patient a priority increase in the deceased-donor (DD) list in the event of another renal failure in the future. This incentive is valuable, considering that a kidney from a living donor typically lasts approximately 15 to 20 years.
@@
The ethical aspects of this policy have been discussed favorably by members of the Canadian transplantation community (Gill et al., 2017). In principle, IKE could be considered for integration into the ongoing reform of the UNOS DD allocation system (Sönmez and Ünver, 2023). However, due to the separate management of living-donor and DD transplants in the U.S., adopting a policy that links the management of these two sources of transplant kidneys would require broader national consensus.
@@
Despite receiving favorable evaluations of its ethical principles within the transplantation community, strictly speaking, IKE falls outside the scope of minimalist market design. While linking the management of living-donor kidneys and DD kidneys may be a valuable policy for increasing the utilization of willing living donors, it is not clear that the absence of this link is the primary cause of the lack of participation by compatible pairs in kidney exchange. In this sense, while IKE offers substantial welfare gains, as an innovation that requires the joint management of multiple separate institutions, a reform based on IKE lacks the typical persuasion strategy inherent in minimalist market design. An ambitious policy proposal for an aspiring design economist, this may be one of the reasons why IKE has fallen short of its practical promise so far, despite the significant welfare gains it offers. [^82] While IKE departs from the idealized version of minimalist market design, it still maintains two of the approach’s three pillars: it identifies the institutional mission, including its ethical considerations, and employs a compelling persuasion strategy. The only pillar IKE does not fully embrace is the avoidance of non-essential changes, as it integrates the management of living-donor and DD kidneys. If separating the management of these kidneys compromises the primary objectives of the transplantation community—a plausible yet unexplored hypothesis—IKE may also satisfy the third pillar over time. In this sense, it is not far from being a minimalist reform. Other policies that involve linking the DD program with living-donor exchange include the list exchange by Ross and Woodle (2000), DD chains by Roth et al. (2004), the voucher program by Veale et al. (2017), and the unpaired kidney exchange by Akbarpour et al. (2024).
7.8 Global Kidney Exchange
Addressing the root causes of institutional failures contributes to the efficacy of minimalist market design. This approach helps avoid inadvertently disrupting essential aspects of an institution while enhancing its performance in other dimensions. Incentivized kidney exchange (IKE) falls short of this feature.
An even more critical factor behind the success of minimalist market design is its alignment with the core objectives of decision makers and stakeholders. These objectives are not only preserved but also fully embraced within IKE. How do policy efforts fare when they disregard this essential feature?
An initiative by Roth and Dr. Mike Rees offers a natural experiment illustrating the importance of this aspect in the effectiveness of economists’ policy work, particularly when conducted as “outsiders” without direct authority. To expand kidney exchange, Roth and Rees promoted the Global Kidney Exchange (GKE) program (Rees et al., 2017, 2022). The program funds kidney transplants for recipients from low- and middle-income countries (LMICs) in exchange for living donors who facilitate transplant cycles or NDD-chains in high-income countries (HICs). One of its goals is to incorporate blood-type-compatible pairs from LMICs, who are unable to afford transplantation, into the kidney exchange pool of HICs.
Since 2015, the Alliance for Paired Kidney Donation (APKD) has implemented GKE (Rees et al., 2022), albeit with limited success. Despite strong advocacy by Roth, a Nobel laureate celebrated for his role in kidney exchange, GKE has resulted in a relatively modest number of transplants. Between January 2015 and February 2022, 17 international patients facing financial barriers to direct transplantation facilitated kidney transplants for 35 U.S. patients through GKE, totaling 52 kidney transplants (Rees et al., 2022). In contrast, over 6,000 kidney exchange transplants occurred in the U.S. during the same period (OPTN, 2024a).
The limited success of GKE, comprising only about 0.5% of kidney exchange transplants in the U.S. from January 2015 to February 2022, can be attributed to strong opposition within the transplantation community. [^83] For a detailed analysis of this opposition, refer to Ambagtsheer et al. (2020). In contrast to our earlier approach in Roth et al. (2003, 2004, 2005b)—which led to NEPKE—critics argue that GKE undermines various ethical norms in transplantation. The persuasion strategies carefully embedded in these initial series of kidney exchange papers—strategies that, in my view, secured our team of economists a major policy victory during our initial interactions with Delmonico—are either bypassed under GKE or have shown limited effectiveness. Indeed, Delmonico, the pivotal figure in the initial collaboration between our team of market designers and medical professionals, has been one of the fiercest opponents of GKE. [^84] Delmonico not only co-authored a comment titled “Opposition to Irresponsible Global Kidney Exchange” in the American Journal of Transplantation (Delmonico and Ascher, 2017) but also spearheaded an ethical objection to GKE through his position as the Executive Director of the Declaration of Istanbul Custodian Group (DICG). Refer to the “Statement of the Declaration of Istanbul Custodian Group Concerning the Ethical Objections to the Proposed Global Kidney Exchange Program” at https://www.declarationofistanbul.org/images/stories/resources/policy_documents/DICG%20Statement%20on%20GKEP_Nov28_final.pdf, last accessed on 11/11/2023.
Opposition to GKE extends beyond the U.S. In 2018, Rees and Roth presented GKE to the National Competent Authorities (NCA) in the European Union. [^85] According to the Directive 2010/53/EU of the European Parliament and of the Council of 7 July 2010 on standards of quality and safety of human organs intended for transplantation, all transplant activities in the EU are overseen by the NCA. Available at https://eur-lex.europa.eu/LexUriServ/LexUriServ.do?uri=OJ:L:2010:207:0014:0029:EN:PDF, last accessed 12/10/2023. In their review, EU NCAs expressed the following ethical concerns, deeming it inappropriate for any EU transplant center to participate in Rees and Roth’s proposed GKE scheme: [^86] See the Statement on Proposed Concept of Global Kidney Exchange by European Union National Competent Authorities on Organ donation and transplantation, available at https://health.ec.europa.eu/system/files/2018-05/nca_statement_gke_adopted_en_0.pdf, last accessed 12/10/2023.
  • GKE has the potential to violate the fundamental principle of non-payment for human organs by providing funding for a kidney transplant procedure to a recipient from a low and middle-income country (LMIC) in exchange for a donor who would facilitate a chain of transplants in a high-income country (HIC).
  • GKE lacks clear and adequate mechanisms to ensure the protection of the LMIC donor, both during the selection process and in the long-term care of the recipient after the transplantation procedure, when they return to their home LMIC.
  • GKE poses a risk of coercion to potential donors in LMIC, who might be offered direct or indirect financial gain in exchange for their kidney.
  • GKE may negatively impact the development of local sustainable donation and transplantation programs in LMIC and initiatives to build ethically sound kidney exchange programs with robust regulatory oversight.
I refrain from taking a position on whether the sharp criticisms of GKE by the EU NCA and other international bodies are justified, as many of these objections fall outside my expertise. Indeed, GKE has also received widespread endorsements, including from prominent bioethicists such as Peter Singer (Minerva et al., 2019) and the Italian government, which recommended GKE to the World Health Organization (Pullen, 2018). What is relevant for my purposes is that GKE faced substantial criticism from various organizations, reducing its effectiveness in achieving the intended welfare gains from kidney exchange. The contrast between the early success of kidney exchange and the sharp opposition to GKE underscores the importance of adopting a cautious approach—a key feature of minimalist market design—when establishing partnerships with experts in other disciplines and policymakers.
@@ 7.9 Kidney Exchange Worldwide: Two Decades On
Attending the 1995 Summer Conference of the Stanford Institute for Theoretical Economics (SITE), just months after receiving my Ph.D., was a pivotal experience that shifted my research focus from abstract theory to use-inspired theory. Eight years later, in August 2003, I presented our kidney exchange research publicly for the first time at another SITE Summer Conference. With its unprecedented potential to save lives through design economics, our work immediately received attention—not only from economists but also from computer scientists.
Within a month, we circulated our findings as an NBER Working Paper Roth et al. (2003) and soon after shared them with Delmonico, a prominent leader in transplantation and the Chief Medical Officer at the New England Organ Bank. At his request, we refined our model in preparation for a potential collaboration.
@@
“This year, more than 60,000 people in the United States will need a kidney transplant. Of that number, roughly 15,000 will receive transplants from either cadavers or living donors. But for those who aren’t so lucky, the prospects have gotten better thanks to a kidney exchange system developed by Harvard economist Alvin Roth, Tayfun Sonmez of Boston College and Utku Unver of University of Pittsburgh. [$\cdots$]
@@
Although this innovation emerged from our NEPKE partnership (Roth et al., 2006), it was the Alliance for Paired Donation (APD) and the National Kidney Registry (NKR) that ultimately implemented and scaled it (Rees et al., 2009). By the early 2010s, this kidney exchange modality had become a cornerstone of the discovery-invention cycle—a series of life-saving innovations that transformed our early theoretical work in on-campus allocation into tangible, real-world healthcare advancements.
@@
7.9.2 The Global Rise of Kidney Exchange Programs Since 2008

### v4

@@ 7.6 Advocating for 3-way Exchanges and NDD-Initiated Chains
To operationalize NEPKE, it was necessary to establish a web-based data entry system. This system would enable each transplant center to enter their patient and donor information directly. The process of setting up this system and addressing bureaucratic requirements for NEPKE spanned about a year. During this period, Ünver developed the software for NEPKE using a priority matching algorithm formulated in Roth et al. (2005b). [^89] A priority matching works by first arranging patient–donor pairs in a priority order and then considering them one by one. At each step, the procedure checks whether the current pair can be part of a feasible outcome while still honoring the transplants promised to higher-priority patients earlier in the process. If it is feasible, the patient in the pair is guaranteed a transplant, though the exact exchange is not finalized until the end if there are multiple possible options. In this way, the final matching ensures that higher-priority patients receive transplants whenever possible. This method produces a Pareto-efficient outcome and, when only two-way exchanges are possible, also maximizes the total number of transplants (Roth et al., 2005b). Even though Delmonico initially indicated that only 2-way kidney exchanges were feasible in New England, for our research purposes, Ünver also coded priority matching algorithms that allowed for larger size exchanges.
After NEPKE became operational in 2006, kidney exchanges in New England were determined by Ünver’s software for a few years. Every few months, our team received anonymized patient–donor data from NEPKE, and we ran the software to derive any new kidney exchanges. This was an exciting period for us, as we were directly impacting and saving lives with our ideas. Hence, upon receiving anonymized data from NEPKE, we would immediately inspect it to estimate how many transplants the software would determine.
Significance of 3-way Exchanges.
In one of these instances, the software produced one more transplant than we thought would be feasible when 3-way exchanges were allowed in the system. This observation came as a big surprise. At that time, our intuition for the maximal-size kidney exchange was derived from Roth et al. (2005b), admittedly in a model that only permitted 2-way exchanges. To address this limitation, we would then utilize some of the technical details in our proofs to estimate the maximum number of transplants if larger-size exchanges were also allowed. This unexpected discovery revealed that our intuition was evidently incomplete. [^90] The maximum number of transplants can be found in polynomial time when only 2-way cycles are allowed, via a reduction to maximum matching (Edmonds, 1965). Once cycles of length three or longer are allowed, however, the problem becomes computationally intractable, as established in earlier work on bounded-length cycle packing (Karp, 1972) and highlighted in the kidney exchange setting by Abraham et al. (2007).
We then conducted a detailed analysis of how allowing larger kidney exchanges affects transplant numbers. This analysis revealed that the underlying partial order structure induced by ABO blood type compatibility leads to two key findings:
@@
    In sufficiently large patient–donor pools, allowing 3-way and 4-way kidney exchanges achieves the maximum number of transplants, and
@@
    Most gains over 2-way exchanges come from 3-way exchanges. [^91] For small pools with 25 patient–donor pairs, the gain was approximately 30% more transplants from kidney exchanges. In large pools with 100 patient–donor pairs, this increase was around 20% more transplants from kidney exchanges.
@@
Given the practical significance of these discoveries, and despite Delmonico’s initial reservations, we convinced our New England partners to include 3-way kidney exchanges in the system, publishing our findings in Saidman et al. (2006) and Roth et al. (2007). [^92] With 3-way exchanges allowed in our system, a priority matching no longer guaranteed the maximization of the number of transplants in the system. Consequently, we adopted an alternative algorithm that maximized the number of transplants à la Roth et al. (2007).
Figure 20: NDD chain with three patient–donor pairs. An altruistic donor donates kidney A to patient P1, initiating the chain. The donor of P1 then donates kidney K1 to P2, whose donor donates K2 to P3. The kidney K3 from the donor of P3 remains available for subsequent transplantation, either to start another chain in the same role as A (effectively extending the current chain over time) or to be allocated to a patient on the deceased-donor list.
Non-Directed Donor (NDD) Donor Chains.
@@
Our collaboration with New England partners also led to another breakthrough. Before this collaboration, altruistic donors were not a major source of living donor kidneys in the U.S. According to the UNOS database, only 8 such transplants were conducted in the U.S. before 2000 (OPTN, 2025). However, this number was increasing in the early 2000s, offering the possibility to initiate donor chains with an altruistic non-directed donor (NDD), rather than a patient–donor pair volunteering for list exchange (see Figure 20 for an illustration). Moreover, since these NDD-chains do not interfere with the DD lists, they were not subject to the same ethical concerns that hindered the utilization of list exchanges in NEPKE.
@@
NDD chains presented another possibility. As mentioned earlier, the Consensus Statement on the Live Organ Donor recommended that all procedures in kidney exchange be performed simultaneously, primarily to prevent a situation in which the donor of a pair gives a kidney as part of an exchange, but their intended patient might not receive a transplant if the other donor backed out or became unavailable. However, this concern was significantly reduced with an NDD chain that starts with an altruistic donation. Because all patients receive a transplant before their donor contributes to the next patient in the chain, the worst-case outcome is simply a shorter chain than its full potential. [^93] Over time, even this worst-case scenario proved to be rare and of negligible consequence. Using data from 344 non-simultaneous NDD chains implemented by the National Kidney Registry from 2008 through May 2016, which generated 1,568 transplants, only 20 of these chains were broken due to reneging (Cowan et al., 2017). Interestingly, the average size of the broken chains was 4.8, compared to 4.5 for unbroken chains.
Based on these observations, we formulated non-simultaneous NDD chains in Roth et al. (2006). [^94] While Roth et al. (2006) was the first to introduce non-simultaneous NDD chains in the literature, a subsequent publication by Lee et al. (2009) revealed that the approach had already been implemented in South Korea. In a multi-center program involving 16 hospitals, 70 such chains were conducted between February 2001 and July 2007, resulting in 179 transplants. While this idea was not pursued by NEPKE, it was later adopted by two other kidney exchange programs in the U.S.: first by the Alliance for Paired Donation (APD), [^95] Starting in 2006, our team also established a partnership with APD. Based on Roth et al. (2006, 2007), APD included 3-way and 4-way exchanges in their allocation mechanism from its inception and incorporated non-simultaneous NDD chains shortly after. Not only did Ünver introduce the concept of non-simultaneous NDD chains to Mike Rees, the head of APD, but the software used by APD from 2006 to 2008–just like the code of NEPKE–was also developed by him. See Anderson et al. (2015) for details of our subsequent partnership with APD, later renamed Alliance for Paired Kidney Donation (APKD). launched in 2006 in the Midwest, and shortly thereafter by the National Kidney Registry, launched in 2008.
@@
Because of its logistical advantages, the role of non-simultaneous NDD chains in U.S. kidney exchange programs grew considerably over time, particularly due to the National Kidney Registry’s success in enrolling altruistic non-directed donors into its system (Cowan et al., 2017; Agarwal et al., 2019).
@@ 7.7 Incentivized Kidney Exchange
While our pioneering role in the evolution and widespread adoption of kidney exchange programs worldwide stands as one of our most consequential policy achievements to date, a crucial policy lever is still missing in a vast majority of these programs—one that has the potential to significantly enhance their efficiency.
@@
As we briefly discussed in Section 7.5, partly due to the circumstances surrounding the foundation of NEPKE—a program that became a blueprint for the rest of the world—participation in kidney exchange programs remains largely restricted to patients who are biologically incompatible with their donors, leading to a substantial efficiency loss.
@@
A handful of policies have been proposed and promoted by various groups to encourage the participation of compatible pairs in kidney exchange programs. Among these, we developed incentivized kidney exchange (IKE) in the early 2010s, as detailed in Sönmez and Ünver (2015) and partially published in Sönmez et al. (2020). Despite its potential, however, this policy has yet to be implemented in practice.
IKE is a targeted strategy for encouraging the participation of blood-type compatible pairs in kidney exchange. The approach hinges on offering patients a higher deceased-donor (DD) priority should they face another renal failure in the future. Since a kidney from a living donor often lasts just 15 to 20 years, this policy provides a practical and forward-looking incentive.
@@
The ethical aspects of this policy have been discussed favorably by members of the Canadian transplantation community (Gill et al., 2017). In principle, IKE could be integrated into the ongoing reform of the UNOS DD allocation system (Sönmez and Ünver, 2023). Yet because living-donor and DD transplants are managed separately in the U.S., adopting a policy that links the two would require broad national consensus.
@@
Despite receiving favorable evaluations of its ethical principles within the transplantation community, strictly speaking, IKE falls outside the scope of minimalist market design. While linking the management of living-donor kidneys and DD kidneys may be a valuable policy for increasing the utilization of willing living donors, it is not clear that the absence of this link is the primary cause of the lack of participation by compatible pairs in kidney exchange. In this sense, while IKE offers substantial welfare gains, as an innovation that requires the joint management of multiple separate institutions, a reform based on IKE lacks the typical persuasion strategy inherent in minimalist market design. As an ambitious policy proposal for an aspiring economic designer, this may be one of the reasons why IKE has fallen short of its practical promise so far, despite the significant welfare gains it offers. [^96] While IKE departs from the idealized version of minimalist market design, it still maintains two of the approach’s three pillars: it identifies the institutional mission, including its ethical considerations, and employs a compelling persuasion strategy. The only pillar IKE does not fully embrace is the avoidance of non-essential changes, as it integrates the management of living-donor and DD kidneys. If separating the management of these kidneys compromises the primary objectives of the transplantation community—a plausible yet unexplored hypothesis—IKE may also satisfy the third pillar over time. In this sense, it is not far from being a minimalist reform. Other policies that involve linking the DD program with living-donor exchange include the list exchange by Ross and Woodle (2000), DD chains by Roth et al. (2004), the voucher program by Veale et al. (2017), and the unpaired kidney exchange by Akbarpour et al. (2024).
@@ 7.8 Kidney Exchange Worldwide: Two Decades On
Attending the 1995 Summer Conference of the Stanford Institute for Theoretical Economics (SITE), just months after receiving my Ph.D., was a pivotal experience that shifted my research focus from abstract theory toward use-inspired theory. Eight years later, in August 2003, I presented our kidney exchange research publicly for the first time at another SITE Summer Conference. With its unprecedented potential to save lives through economic design, our work immediately drew attention from economists, computer scientists, and operations researchers alike.
Within a month, we circulated our findings as an NBER Working Paper (Roth et al., 2003) and soon after shared them with Delmonico, a prominent leader in transplantation and the Chief Medical Officer at the New England Organ Bank. At his request, we refined our model in preparation for a potential collaboration.
@@
“This year, more than 60,000 people in the United States will need a kidney transplant. Of that number, roughly 15,000 will receive transplants from either cadavers or living donors. But for those who aren’t so lucky, the prospects have gotten better thanks to a kidney exchange system developed by Harvard economist Alvin Roth, Tayfun Sonmez of Boston College and Utku Unver of University of Pittsburgh. $\cdots$]
@@
Although this innovation emerged from our NEPKE partnership (Roth et al., 2006), it was the Alliance for Paired Donation (APD) and the National Kidney Registry (NKR) that ultimately implemented and scaled it (Rees et al., 2009). By the early 2010s, as depicted in Figure 37, this kidney exchange modality had become a cornerstone of the discovery–invention cycle—a series of life-saving innovations that transformed our early theoretical work on housing allocation into tangible, real-world healthcare advancements.
@@
7.8.2 The Global Expansion of Kidney Exchange Since the Mid-2000s